Pretend you recently attended a wedding reception. After the event, 90 of the 120 attendees reported having severe gastroenteritis for the weekend. The epidemiologist in you wanted to figure out which food(s) at the event might be responsible, so you interviewed all of the attendees. Of 60 people who recalled eating mini beef Wellingtons, 50 reported getting sick. Of 100 people who recalled eating spicy chicken satay, 85 reported getting sick. Of 40 people who recalled eating vegetables with ranch dressing dip, 35 reported getting sick. Of 75 people who recalled eating shrimp kabobs with cocktail sauce, 58 reported getting sick.

a) Indicate how the components of an epidemiologic triangle could be implicated in this illness.

b) What is the mode of transmission, portal of entry and portal of exit for the possible infectious agent in this illness?

c) Calculate the attack rates of gastroenteritis for each food involved.
Answer #1

a) Gastroenteritis defines as inflammation of stomach and intestines as a result if exposure to bacteria or viral factors. Epidemiological triad is most accurate model describing disease causation. The triangle consists of three components that includes an external agent,susceptible host and a favorable environment. This models helps to understand the cause and effect relationship. In this scenario as per the epidemiological component agent is the eating of food either not cooked properly or contaminated with organisms. Host is the human being who consumed the contaminated food. Environment factors are the situation where the outbreak of disease happened, here the wedding reception is the environment where many people gathered and consumed food, the quality of food create impact of illness.

b) Mode of transmission refers to the way how the infectious organisms moves from one host to the other. The modes of transmission includes direct transmission from one person to other consumption of contaminated food and water. Portal of entry is the way the organisms attacks or enters into the host system. Ingestion of contaminated food that enters gastrointestinal system and initiates infection. Portal of exit defines as a site from where the microorganisms leave the host. Gastroenteritis the portal of entry is through vomitus or feces.

c) Attack rate is a tool used in epidemiological studies to calculate the frequency of disease and the population at the risk for disease. The formula to calculate attack rate are: Number of people exposed who are sick divided by total number of people exposed.

Of 60 people who recalled eating mini beef Wellingtons, 50 reported getting sick.

Attack rate = 50 / 60 x 100 = 83%

Of 100 people who recalled eating spicy chicken satay, 85 reported getting sick.

Attack rate = 85 / 100 x 100 = 85%

Of 40 people who recalled eating vegetables with ranch dressing dip, 35 reported getting sick.

Attack rate = 35 / 40 x 100 = 88%

Of 75 people who recalled eating shrimp kabobs with cocktail sauce, 58 reported getting sick.

Attack rate = 58 / 75 x 100 = 77%